OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? x − 3y = 9 x + 3y = 9 3x − y = −9 3x + y = 9

jimthompson5910 (jim_thompson5910):

Since you're given the answers, you can check them x − 3y = 9 3 − 3(2) = 9 ... plug in x = 3 and y = 2 (from the point (3,2)) 3 - 6 = 9 -3 = 9 which is false So (3,2) does NOT lie on the line x-3y = 9

jimthompson5910 (jim_thompson5910):

check all the possible answer choices with the given points

jimthompson5910 (jim_thompson5910):

The line where both points work will be your answer

OpenStudy (anonymous):

Isn't that x + 3y = 9? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you are absolutely correct

OpenStudy (anonymous):

Thank you! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw